### v1

@@ 2.3.1 Proposal 1: Definition of support and diagnostic using High-Density Region (HDR)
This metric captures the extent to which a specific intervention is supported within the population. It represents the proportion of individuals for whom the confounder distributions under the intervention do not sufficiently overlap, indicating scenarios where the intervention is either highly improbable or practically infeasible for certain individuals.
The non-overlap ratio ranges from 00 to $1$:
  • A value near 00 signifies strong support for the positivity assumption at this intervention value, indicating that almost no positivity violation occurs at the current coverage level.
  • A value close to $1$ indicates severe positivity violations, where the intervention $a$ is largely unsupported by the data. This lack of overlap suggests that reliable causal inferences for $a$ may be unattainable due to insufficient data support.
By identifying regions of the intervention space with inadequate overlap, the non-overlap ratio serves as a critical diagnostic tool for evaluating the validity of causal assumptions. Its practical application will be detailed in the subsequent sections.
3 Novel Estimands to Address Positivity Violations
3.1 Standard Causal Dose-Response-Curve
{boxwithhead}
Estimand 1 (Standard) The standard causal dose-response curve is a function $m^{\text{standard}}:\mathcal{A}\to\mathcal{Y};\ a\mapsto E(Y^{a})$, which can be identified as
$m^{\text{standard}}(a)$ =$\overset{(i)}{=}E[E(Y^{a}|A=a,\mathbf{L})]\overset{(ii)}{=}E[E(Y|A=a,\mathbf{L})]$ (6)
[…]
where equality $(i)$ follows by the law of iterated expectation and conditional exchangeability, and equality $(ii)$ is due to the consistency and positivity assumption (See Equation 1).
The standard causal dose-response curve, reliant on the positivity assumption in Equation (1), poses significant challenges due to frequent positivity violations, especially with continuous interventions. As the range of an intervention widens, certain regions of the confounder space often lack sufficient support, leading to extrapolation and unreliable estimates.
3.2 Causal Dose-Response-Curve with Data-Adaptive Intervention
To address these challenges, we propose a novel estimand that mitigates the impact of positivity violations by focusing on well-supported regions of the intervention space, guided by the diagnostic tool proposed in previous section. This approach can effectively relax the positivity assumption.
3.2.1 Data-Adaptive Intervention
Unlike static interventions, which assign a fixed treatment $a$ to all units, or dynamic interventions, which vary based on individual confounders, we introduce a more refined approach called the data-adaptive intervention. This strategy customizes the intervention for each individual based on their confounders and the observed data.
For a given oracle intervention value $a$, the concept of a data-adaptive intervention can be outlined as follows:
$\begin{cases}\text{Intervene with $a$,}&\text{if it is supported in the data},\\ \text{Intervene adaptively,}&\text{otherwise}.\end{cases}$ (7)
Formally, the data-adaptive intervention $\mathbb{d}$ is a data-adaptive function that maps an intervention value $a$, confounders $\mathbf{l}$ and the observed data $\mathcal{D}$ into a new intervention value, which is defined as:
$\mathbb{d}(a,\mathbf{l};\mathcal{D})=\begin{cases}a,&\text{if }a\in\mathcal{A}_{\alpha}(\mathbf{l}),\\ h(a,\mathbf{l};\mathcal{D}),&\text{otherwise}.\end{cases}$ (8)
$\mathcal{A}_{\alpha}(\mathbf{l})$ is the high-density region defined in Section 2.3. The function $h$ is a data-adaptive function selecting a substitution intervention based on the intervention value $a$, confounders $\mathbf{l}$ and data $\mathcal{D}$.
3.2.2 High Density Region
In order to determine whether an intervention is supported by the data, we adopt here the high-density regions introduced in Section 2.3 as a support indicator.
@@ 3.2.3 The Most Feasible Intervention
A natural decision for $h$ is to substitute the intervention when it is not supported given the confounders, while ensuring the substitution is as close as possible to the oracle intervention:
$h(a,\mathbf{l};\mathcal{D})=\underset{a^{*}:a^{*}\in\mathcal{A}_{\alpha}(\mathbf{l})}{\mathrm{argmin}}|a^{*}-a|$ (9)
An illustration is provided in Figure 2, where the blue band represents the feasible region for the intervention given the confounders. The most feasible intervention is the closest supported intervention to the oracle intervention of interest, shown in yellow crosses.
Figure 2: Illustration of the most feasible intervention in Eqn. 10
3.2.4 Proposal 2: Fix the positivity issue detected in Proposal 1 through novel estimand
{boxwithhead}
Estimand 2 (Feasible) With both definitions of support and substitution intervention function, in Equation (4) and (9), our proposed new estimand is given by:
$m^{\text{feasible}}:a\mapsto E(Y^{a^{\mathbb{d}}}),$ (10)
where $a^{\mathbb{d}}:=\mathbb{d}(a,\mathbf{l};\mathcal{D})$ with $\mathbb{d}$ defined in Equation (8) and $h$ defined in Equation (9).
This can also be written as
$m^{\text{feasible}}:a\mapsto$ $P(a\in\mathcal{A}_{\alpha})\cdot E(Y^{a})+$ (11)
$(1-P(a\in\mathcal{A}_{\alpha}))\cdot E(Y^{a^{\mathbb{d}}})$
@@ 3.2.4 Proposal 2: Fix the positivity issue detected in Proposal 1 through novel estimand
Figure 2 illustrates the concept of the most feasible intervention and its interpretation using pseudo-data samples. The dots, shown in red or grey, represent observed data points. We begin by considering an ideal or oracle intervention target $A=a$, represented by the orange horizontal line, which is intended to evaluate a causal quantity such as the standard causal dose-response curve (CDRC) at this target intervention value $a$. This target quantity, which counterfactually assigns the fixed intervention value $a$ to each subject, assumes positivity. However, while this oracle intervention is theoretically appealing, it may not always be supported by sufficient data across all confounder combinations. The shaded blue area in the figure represents the high-density region, where data is abundant, capturing the most likely combinations of confounders and interventions. The observed data points are colored black if the target intervention $a$ is within their conditional support, and grey otherwise. The orange crosses indicate the most feasible intervention of each subject for intervention target $a$. If the target intervention $a$ is supported, the feasible intervention coincides with $a$; otherwise, it corresponds to the closest intervention value to $a$ that lies within the high-density region.
@@
Violations of the positivity assumption are not merely theoretical concerns but also reflects real-world scenarios where certain interventions are impractical or unrealistic for specific subpopulations. For example, consider drug concentration levels across patients, as discussed in Section 5. Patients with a fast metabolism, who metabolize drugs rapidly, naturally exhibit low drug concentrations even when administered standard doses. For these patients, a high drug concentration target would be unrealistic because it would require doses far beyond what is commonly observed or safely prescribed. While theoretically possible, achieving such concentrations would involve dosing regimens that are rarely practiced due to clinical and practical concerns. This leads to a situation where data corresponding to such high concentrations is either extremely sparse or non-existent, leading to biased causal inferences due to the lack of overlap and making causal conclusions drawn from this target intervention more likely to be unreliable.
@@
The proposed CDRC with the most feasible intervention addresses these issues by adapting the intervention target to stay within regions where data is enough and well-supported. Instead of forcing an unrealistic target that fails to align with real-world conditions, this approach adjusts the intervention based on what is practically achievable and observable within the data. This has two key benefits: first, it reduces bias caused by data sparsity and positivity violations, and second, it ensures that the causal effects being estimated are relevant and actionable in practice.
@@
In contrast to fixed interventions, which risk oversimplifying or misrepresenting causal relationships by imposing uniform targets, the most feasible intervention approach offers a more nuanced and reasonable strategy. It acknowledges the inherent variability in real-world data and adapts the intervention accordingly, leading to more robust, reliable, and meaningful causal inferences that better reflect both the data and the practical realities of clinical and observational settings.
3.3 Trimming
In causal estimation, trimming refers to the process of excluding certain observations from the analysis to improve the accuracy and robustness of estimated treatment effects 6. This is typically done by removing observations that have a low conditional density given the confounders, which can help reduce the impact of outliers or extreme values that may introduce bias or distort the results as given in Equation (12).
$m^{\text{trimming}}(a):a\mapsto E(Y^{a}|f(a|\mathbf{L})>c),$ (12)
where $c$ is a critical threshold for the conditional density of the intervention given the confounders.
{boxwithhead}
Estimand 3 (Trimming) The trimming estimand can also be based on the non-overlap ratio $\tau$ defined in Equation (5):
$m^{\text{trimming}}(a):a\mapsto E(Y^{a}|a\in\mathcal{A}_{\alpha}(\mathbf{l}))$ (13)
which can be identified by
[…]
where $\text{trimming}(a,\mathbf{l})$ is defined as:
$\text{trimming}(a,\mathbf{l})=\begin{cases}\dfrac{1}{\tau},&\text{if }a\notin\mathcal{A}_{\alpha}(\mathbf{l}),\\ 0,&\text{otherwise}.\end{cases}$ (15)
Here, $\mathcal{A}_{\alpha}(\mathbf{l})$ and $\tau$ are defined as in Section 2.3.
@@ 3.4.1 Causal Estimation
To estimate the standard CDRC in (6), parametric g-formula can be easily applied for the single time-point intervention scenario. Substitution estimators can be constructed to estimate (10) with detailed procedure outlined in Algorithm 1.
Algorithm 1 Estimation Algorithm for $m^{\text{standard}}$, $m^{\text{feasible}}$, $m^{\text{trimming}}$ and the Non-Overlap Ratio $\tau$
Step 1: Set-Up
Let $j$ represent the number of subjects in the dataset. Define a set of interventions of interest $\overline{\mathbf{a}}=(a_{1},\cdots,a_{m})^{\mathrm{T}}$, which corresponds to the potential treatments or actions under analysis. Define the coverage level $\alpha$ to identify the high-density regions.
Step 2: Identify High-Density Regions
  • (a) Conditional Density Estimation: Estimate the conditional density model $\hat{f}(a|\mathbf{l})$ using the observed data $\mathcal{D}$.
  • (b) High-Density Threshold: For each subject $j$, determine the high-density threshold $\hat{f}^{j}_{\alpha}$ by computing the $\alpha$-quantile of $\hat{f}(a|\mathbf{l}_{j})$.
Step 3: Compute the Non-Overlap Ratio
For each intervention $a_{i}$, calculate the non-overlap ratio as:
$\hat{\tau}(a_{i})=\frac{1}{n}\sum_{j=1}^{n}I(\hat{f}(a_{i}|\mathbf{l}_{j})<\hat{f}_{\alpha}^{j})$
Step 4: Outcome Model Estimation
Estimate the conditional expectation model $\hat{E}(Y|A,\mathbf{L})$.
Step 5: Estimate Expected Outcomes Using Plug-in Estimators
  • (a) Estimand 1 (standard): For each intervention $a_{i}$, calculate the standard estimand:
    $\hat{m}^{\text{standard}}(a_{i})=\frac{1}{n}\sum_{j=1}^{n}\hat{E}(Y|A=a_{i},\mathbf{l}_{j})$
  • (b) Estimand 2 (feasible): For each intervention $a_{i}$, for each subject $j$, evaluate the conditional density as estimated in Step 2(a). Estimate $\hat{\mathbb{d}}(a_{i},\mathbf{l}_{j};\mathcal{D})$ as follows: if $\hat{f}(a_{i}|\mathbf{l}_{j})$ is below the high-density threshold $\hat{f}_{\alpha}^{j}$, adjust $a_{i}$ to the nearest feasible value in $\overline{\mathbf{a}}$; otherwise, retain $a_{i}$. Then, calculate the feasible estimand:
    $\hat{m}^{\text{feasible}}(a_{i})=\frac{1}{n}\sum_{j=1}^{n}\hat{E}(Y|A=\hat{\mathbb{d}}(a_{i},\mathbf{l}_{j};\mathcal{D}),\mathbf{l}_{j})$
  • (c) Estimand 3 (trimming): For each intervention $a_{i}$, for each subject $j$, evaluate the conditional density as estimated in Step 2(a). Estimate the trimming function $\hat{\mathrm{trimming}}(a_{i},\mathbf{l}_{j})$ as:
    $\hat{\mathrm{trimming}}(a_{i},\mathbf{l}_{j})=\begin{cases}\dfrac{1}{\hat{\tau}(a_{i})},&\text{if }\hat{f}(a_{i}|\mathbf{l}_{j})\geqslant\hat{f}_{\alpha}^{j},\\ 0,&\text{otherwise}.\end{cases}$
    Then, calculate the trimming estimand:
    $\hat{m}^{\text{trimming}}(a_{i})=\frac{1}{n}\sum_{j=1}^{n}\hat{E}(Y|A=a_{i},\mathbf{l}_{j})\cdot\hat{\mathrm{trimming}}(a_{i},\mathbf{l}_{j})$
Step 6: Estimate Compatibility Intervals via Bootstrapping
Use bootstrapping to calculate compatibility intervals (CIs) for $\hat{m}^{\text{standard}}$, $\hat{m}^{\text{feasible}}$, and $\hat{m}^{\text{trimming}}$.

### v2

@@ 3.3.1 Data-Adaptive Intervention
Formally, let $\mathcal{A}_{\alpha}(\mathbf{l};P)$ denote the highest density regions (HDRs) of treatment values supported by the distribution $P$ given covariates $\mathbf{l}$, as defined in Section 2.2.
We define the data-adaptive intervention operator
$\mathop{}\!\mathrm{d}:\mathcal{A}\times\mathcal{L}\times\mathcal{P}\rightarrow\mathcal{A},$
where $\mathcal{P}$ denotes the space of candidate distributions. The operator assigns to each $(a,\mathbf{l},P)$ a feasible treatment value via
$\mathop{}\!\mathrm{d}(a,\mathbf{l};P)=\begin{cases}a,&\text{if }a\in\mathcal{A}_{\alpha}(\mathbf{l};P),\\[6.0pt] h(a,\mathbf{l};P),&\text{otherwise},\end{cases}$ (11)
where $h:\mathcal{A}\times\mathcal{L}\times\mathcal{P}\to\mathcal{A}$ is a substitution function satisfying
$h(a,\mathbf{l};P)\in\mathcal{A}_{\alpha}(\mathbf{l};P)\quad\text{whenever }a\notin\mathcal{A}_{\alpha}(\mathbf{l};P).$ (12)
The corresponding counterfactual outcome under the data-adaptive intervention is denoted
$Y^{\mathop{}\!\mathrm{d}}\equiv Y^{\mathop{}\!\mathrm{d}(a,\mathbf{l};P)}.$
In practice, the true distribution $P$ is unknown and replaced by the empirical distribution $P_{n}$, which makes the rule inherently data-adaptive. For notational convenience, write
$\mathcal{A}_{\alpha}(\mathbf{l}):=\mathcal{A}_{\alpha}(\mathbf{l};P_{n})\quad\text{and}\quad h(a,\mathbf{l}):=h(a,\mathbf{l};P_{n}).$
Then the data-adaptive intervention rule becomes
$\mathop{}\!\mathrm{d}(a,\mathbf{l};P_{n})=\begin{cases}a,&\text{if }a\in\mathcal{A}_{\alpha}(\mathbf{l}),\\[6.0pt] h(a,\mathbf{l}),&\text{otherwise}.\end{cases}$ (13)
@@ 3.3.2 The Most Feasible Intervention
A natural choice for the substitution function $h$ is to reassign unsupported interventions to the nearest feasible value, ensuring minimal deviation from the intended intervention:
$h(a,\mathbf{l})=\underset{a^{*}\in\mathcal{A}_{\alpha}(\mathbf{l})}{\arg\min}\,|a^{*}-a|.$ (14)
An illustration is provided in Figure 2a. The horizontal axis shows intervention values and the vertical axis their estimated conditional density. Green shaded areas mark the highest density regions (HDRs), where the conditional density exceeds the threshold $f_{\alpha}$ (dashed line) and interventions are well supported; grey areas indicate poorly supported regions. The red arrows denote hypothetical targets: Target 1 lies outside the HDR, while Target 2 lies inside. Under the proposed rule, unsupported targets are reassigned to the nearest feasible value within the HDRs (green arrows). Thus, Target 1 shifts to the closest HDR point, while Target 2 remains unchanged. This illustrates the principle of the most feasible intervention: retain the target if supported, otherwise reassign it to the nearest supported value.
(a) Targets outside the HDRs are reassigned to the closest feasible value within the HDRs. The green shaded area represents the highest density region (well-supported interventions), the grey area indicates regions with little or no support, and the dashed line marks the density threshold $f_{\alpha}$. Red arrows denote target interventions: Target 1 lies outside the HDR and is reassigned to the nearest feasible value (Most Feasible Intervention 1), while Target 2 lies within the HDR and is retained unchanged (Most Feasible Intervention 2).
(b) Targets inside the HDRs remain unchanged, while unsupported targets are shifted adaptively. Red circles denote target interventions, green circles represent their most feasible reassigned values, and black squares indicate observed data points across strata of the confounders. Green shaded areas indicate HDRs, while grey areas mark unsupported regions. This illustrates how infeasible interventions are reassigned adaptively to the closest feasible value, ensuring that interventions are only minimally adjusted.
Figure 2: Illustration of the most feasible intervention defined in Equations 13 and 14. Panel (a) shows the conditional density of treatment values given covariates, with HDRs (green) representing well-supported regions and grey areas denoting poorly supported regions. Target interventions (red arrows) are either retained if they fall inside HDRs or reassigned to the nearest feasible value within HDRs (green arrows). Panel (b) extends this principle across confounder strata, where targets (red circles) are reassigned to the closest supported value (green circles) whenever they fall outside HDRs, while feasible targets remain unchanged. Together, the panels illustrate the principle of the most feasible intervention: retain the target when supported, and otherwise make the smallest necessary adjustment to ensure feasibility.
3.3.3 Proposal 2: A Novel Estimand Based on Feasible Interventions
Estimand 3 (Feasible) With the definitions of support and the substitution function in Equations 4 and 14, the proposed estimand is $m^{\text{feasible}}(a)=\mathbb{E}\!\left(Y^{\mathop{}\!\mathrm{d}}\right),$ (15) where $\mathop{}\!\mathrm{d}$ is defined in Equation 13 and $h$ in Equation 14. Equivalently, $m^{\text{feasible}}(a)=\mathbb{E}\!\Big[Y^{a}\mathbbm{1}\{a\in\mathcal{A}_{\alpha}(\mathbf{L})\}+Y^{\,h(a,\mathbf{L})}\mathbbm{1}\{a\notin\mathcal{A}_{\alpha}(\mathbf{L})\}\Big].$ (16)
In words, each unit either (i) receives the target intervention $a$ if $a$ lies in its supported region $\mathcal{A}_{\alpha}(\mathbf{L})$, or (ii) is reassigned to the nearest feasible value $h(a,\mathbf{L})$ otherwise. The resulting curve averages these counterfactuals across the population, yielding a causal dose–response curve that closely aligns with the original estimand while remaining identifiable in the presence of positivity violations. By limiting substitutions to the minimal adjustment necessary, this approach preserves interpretability and avoids the distortions introduced by extrapolation into unsupported regions.
Comparison of Estimand 2 and Estimand 3
Both the weighted CDRC and the feasible CDRC aim to address positivity violations by restricting attention to regions with sufficient empirical support. However, they adapt the estimand in fundamentally different ways. The weighted CDRC of Schomaker et al. [31] applies a fixed cutoff $c$ to the conditional density and redefines the target through projection weights. This ensures identifiability without requiring extrapolation, but it alters the causal parameter: in regions of support, the curve coincides with the standard CDRC, while in unsupported strata it relies on weighted associational components. As Schomaker et al. emphasize, this means the weighted curve sticks to the causal question where feasible but falls back on associations where necessary, yielding a compromise between interpretability and stability. By contrast, the feasible CDRC employs a unit-specific, data-adaptive threshold and modifies intervention assignments explicitly by reassigning unsupported values to the closest feasible ones. This produces a transparent intervention rule targeting the well-defined quantity $E(Y^{\mathop{}\!\mathrm{d}})$. In short, the weighted CDRC trades some causal interpretability for broader identifiability, whereas the feasible CDRC preserves interpretability while adapting directly to the observed support.
@@ 3.3.3 Proposal 2: A Novel Estimand Based on Feasible Interventions
Figure 2b illustrates how the most feasible intervention strategy operates at the individual level, conditional on confounders. The vertical axis represents confounder values, and the horizontal axis represents the intervention. The shaded green band marks the highest density regions, where intervention values are well supported by the data, while the surrounding grey area corresponds to lowest density regions with little or no support. The red vertical line indicates the oracle intervention target, which may not be feasible for all individuals. Units whose oracle intervention lies outside the highest density regions (shown as red circles) cannot be assigned this target directly. Instead, the proposed rule reassigns them to the nearest feasible value within the highest density regions, shown by the green circles and arrows. By contrast, units whose oracle target already falls within the highest density regions (green circles on the red line) retain their original assignment. Black squares represent the observed data (intervention–confounder pairs), from which the highest density regions are estimated. Overall, the figure demonstrates that the most feasible intervention rule systematically reassigns unsupported targets to the closest feasible values, ensuring that every individual receives an intervention assignment compatible with the observed data distribution.
@@
Violations of the positivity assumption are not merely theoretical concerns but also reflect real-world scenarios where certain interventions are impractical or unrealistic for specific subpopulations. For example, consider drug concentration levels across patients, as discussed in Section 5. Patients with a fast metabolism, who metabolize drugs rapidly, naturally exhibit low drug concentrations even when administered standard doses. For these patients, a high drug concentration target would be unrealistic because it would require doses far beyond what is commonly observed or safely prescribed. While theoretically possible, achieving such concentrations would involve dosing regimens that are rarely practiced due to clinical and practical concerns. This leads to a situation where data corresponding to such high concentrations is either extremely sparse or non-existent, resulting in biased causal inferences and making conclusions drawn from such targets unreliable.
@@
The proposed CDRC with the most feasible intervention addresses these issues by adapting the intervention target to stay within regions where data are sufficiently abundant and well-supported. Instead of forcing an unrealistic target that fails to align with real-world conditions, this approach adjusts the intervention based on what is practically achievable and observable within the data. This has two key benefits: first, it reduces bias caused by data sparsity and positivity violations, and second, it ensures that the causal effects being estimated are relevant and actionable in practice.
@@
Unlike fixed interventions, which risk oversimplifying causal relationships by imposing uniform targets, the most feasible intervention approach offers a more nuanced and reasonable strategy. It acknowledges the inherent variability in real-world data and adapts the intervention accordingly, leading to more robust, reliable, and meaningful causal inferences that better reflect both the data and the practical realities of clinical and observational settings.
3.3.4 Data-Adaptive Trimming Causal Dose–Response Curve
In causal estimation, trimming refers to the practice of excluding observations with insufficient treatment support in order to stabilize estimation and mitigate the impact of positivity violations [7]. In its classical form, trimming relies on a fixed threshold $c>0$ applied to the conditional treatment density, yielding
$m^{\text{trimming}}(a):a\mapsto\mathbb{E}\!\left(Y^{a}\,\big|\,f_{A|\mathbf{L}}(a\mid\mathbf{L})>c\right),$ (17)
where $f_{A|\mathbf{L}}$ is the conditional density of $A$ given $\mathbf{L}$.
Estimand 4 (Trimming) In our framework, trimming is expressed in a more flexible, data-adaptive way using the empirical feasible set $\mathcal{A}_{\alpha}(\mathbf{L})$ introduced in Section 2.2. Specifically, we define $m^{\text{trimming}}(a):a\mapsto\mathbb{E}\!\left(Y^{a}\,\big|\,a\in\mathcal{A}_{\alpha}(\mathbf{L})\right),$ (18) that is, the mean counterfactual outcome restricted to those units for whom the intervention value $a$ lies within the empirically supported region given their covariates. This estimand can be identified by $\int\mathbb{E}(Y\mid A=a,\mathbf{L})\;t(a,\mathbf{L})\,dP_{\mathbf{L}}(\mathbf{l}),$ (19) where the trimming weight is $t(a,\mathbf{l})=\begin{cases}\dfrac{1}{\tau},&\text{if }a\in\mathcal{A}_{\alpha}(\mathbf{L}),\\[6.0pt] 0,&\text{otherwise},\end{cases}$ (20) with $\tau=P_{0}\!\big(a\in\mathcal{A}_{\alpha}(\mathbf{L})\big)$ denoting the probability that $a$ falls within the empirical feasible region.
Comparison: This data-adaptive trimming estimand improves upon classical trimming by replacing an arbitrary, fixed threshold with a threshold derived from the observed distribution of the data. In doing so, it ensures that only well-supported units contribute to estimation, while retaining interpretability by conditioning on the feasible region. However, like classical trimming, it targets a restricted population rather than the full sample, which may reduce generalizability. In contrast, the feasible CDRC (Estimand 3) reassigns unsupported interventions rather than excluding them, preserving the entire study population while still avoiding extrapolation.
@@ 3.4.1 Causal Estimation
To estimate the standard CDRC in Equation (6), a parametric g-formula can be readily applied for single time-point interventions. Substitution (plug-in) estimators can also be constructed to estimate the data-adaptive estimands introduced in Equation (15) and (18), with the full procedure detailed in Algorithm 1.
Notes:
  1. 1)
    The estimated conditional treatment density $\hat{f}(a\mid\mathbf{l})$ (Step 2) is used only to define the data-adaptive estimands (feasible and trimming) and does not enter into the estimation of the outcome model.
  2. 2)
    Valid statistical inference using bootstrap intervals (Step 6) requires, under standard regularity, estimators that are $\sqrt{n}$-consistent and asymptotically linear with a well-defined influence function[3, 34, 35, 23]. These conditions are satisfied, for example, by correctly specified parametric outcome models, and can also be met by certain flexible estimators such as the Highly Adaptive Lasso (HAL) [2], which under mild conditions yields plug-in estimators that are $\sqrt{n}$-consistent and asymptotically linear in pathwise differentiable settings [6, 33]. The CDRC, however, is not pathwise differentiable as a full curve, so bootstrap validity cannot be guaranteed in general. At best, bootstrap may provide pointwise inference at fixed treatment values under correct parametric specification. In contrast, HAL plug-in estimators of the CDRC are pointwise asymptotically normal under undersmoothing, which enables valid delta-method inference based on the influence curve of the HAL coefficients [32]. This approach is analogous to the empirical sandwich variance estimator recently proposed for ICE g-computation by Zivich et al. [39].
  3. 3)
    When the target estimand is defined in a data-adaptive manner—such as via estimated highest density regions—standard independence assumptions are violated. As emphasized by Hubbard et al.[19], the intervention assigned to a given unit depends on functions of the data estimated from other units, introducing dependence between the observed data and the target parameter. This dependence may affect the validity of standard inferential procedures such as the bootstrap.
  4. 4)
    A proposed remedy is sample splitting or cross-validation, whereby the data-adaptive rule is defined on one subset of the data and the parameter is estimated on another. This restores approximate independence between the estimand and the estimation process. However, such approaches introduce practical complexity: they yield multiple, split-specific estimands, and averaging across them may reduce interpretability.
  5. 5)
    In this work, we choose to use the full sample to define and estimate the data-adaptive estimands. This preserves a single, coherent target parameter that reflects the region of intervention support within the observed data. While this choice introduces some additional dependence between the data and the estimand, we view it as a pragmatic balance between interpretability and statistical rigor.
  6. 6)
    Our simulation study (Appendix D) includes an empirical evaluation of bootstrap coverage, both with and without sample splitting. In the settings considered, bootstrap intervals achieve coverage close to nominal levels in well-supported regions. Thus, the practical benefit in interpretability outweighs the modest loss of conservativeness in inference for many applied scenarios.